OpenStudy (anonymous):

Which function is the inverse of f(x) = 1/4 x^3 + 1 A) F^-1(x)= 3 SR 4x B) F^-1(x)=3 SR x-1 C) F^-1(x)= 3 SR 4x-4 D) F^-1(x)= 3 SR 4x-1

OpenStudy (anonymous):

I have no idea what to do here. So if anyone could guide me through it I would greatly appreciate it.

OpenStudy (anonymous):

let f(x) = y first.. So: \[\large y = \frac{1}{4}x^3 + 1\] Right??

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

You must do the inverse operations in inverse order. This equation is x cubed, then times 4 then plus 1. So the inverse is: ...

OpenStudy (anonymous):

So: \[y - 1 = \frac{1}{4} x^3\] \[x^3 = 4y - 4 \implies x = \sqrt[3]{4y - 4}\] Just swap x and y here

OpenStudy (anonymous):

replace x by y and y with x here you will get \(f^{-1}(x)\)..

OpenStudy (anonymous):

Welcome dear...